# 2cm2cm1cm1cm

### v1

@@ 3.3.1 Unbound pair scattering channel
Far apart from each other, two fermions only experience a very weak repulsion force due to the extended Hubbard interaction while the probability that they bind together to form a boson is also very small. Thus, in this situation, one expects that the dynamics of such a pair is governed by the fermionic part, and even by the hopping term alone. During intermediate times they could of course interact, as they may get close to each other, and they could even be bound together via the effective attraction caused by fermion-boson exchange processes. The lifetime of bound fermions should however be finite in this situation and they are expected to be released at some point and behave again as two free fermions that go far apart from each other for large times. We show below that this heuristics can be put in precise mathematical terms.
@@
$M_{\mathfrak{f}\left(k\right)}$ being the fiber Hamiltonian defined as the multiplication operator by $\mathfrak{f}(k)\in C(\mathbb{T}^{2})$ (see (28)) while $P_{x}$ is the orthogonal projection onto the one-dimensional subspace $\mathbb{C}\mathfrak{\hat{e}}_{x}\subseteq L^{2}(\mathbb{T}^{2})$. Compare (52) with Equations (40)–(31) and Proposition 2.1. By Lemma 4.20, note that $P_{\mathrm{ac}}(\mathrm{H}_{f})=\mathfrak{1}$.
@@
See Equation (23). Note that $\mathfrak{U}$ is an isometry, i.e., a norm preserving linear transformation. In fact, it is the canonical fiberwise inclusion of $\mathfrak{H}_{f}$ into $L^{2}\left(\mathbb{T}^{2},\mathcal{H}\right)$. Recall from Proposition 2.1 that $A\left(\mathrm{U},\cdot\right)$, defined by (31)–(40), is the fiber decomposition of the operator $UHU^{\ast}$. Then, we obtain wave and scattering operators with respect to fermionic parts:
@@
    Wave operators (see (48)):
    $W^{\pm}\left(UHU^{\ast},\mathrm{H}_{f};\mathfrak{U}\right)=\left({\int_{\mathbb{T}^{2}}^{\oplus}}W^{\pm}\left(A\left(k\right),\left(M_{\mathfrak{f}\left(k\right)}+R\left(\mathrm{V},\mathrm{v}\right)\right)\oplus A_{2,2}\left(k\right)\right)\nu\left(\mathrm{d}k\right)\right)\mathfrak{U}$
@@
    $\mathrm{ran}\left(W^{\pm}\left(UHU^{\ast},\mathrm{H}_{f};\mathfrak{U}\right)\right)=\int_{\mathbb{T}^{2}}^{\oplus}L^{2}\left(\mathbb{T}^{2}\right)\oplus\{0\}\,\nu\left(\mathrm{d}k\right)\ .$
@@
    Scattering operator (see (50)):
    $S\left(UHU^{\ast},\mathrm{H}_{f};\mathfrak{U}\right)=\mathfrak{U}^{\ast}\left({\int_{\mathbb{T}^{2}}^{\oplus}}S\left(A\left(k\right),\left(M_{\mathfrak{f}\left(k\right)}+R\left(\mathrm{V},\mathrm{v}\right)\right)\oplus A_{2,2}\left(k\right)\right)\nu\left(\mathrm{d}k\right)\right)\mathfrak{U}\ .$
Proof. Observe that the operator difference $(UHU^{\ast}-\mathrm{H}_{f})$ is not trace-class (it is not even compact) and, thus, the existence of this scattering channel is not a direct consequence of the well-known Kato-Rosenblum theorem [35, Theorem XI.8]. In fact, one of the main steps of the proof is to show that this difference is the direct integral of a strongly measurable family of trace-class operators. By this means we are then able to apply the Kato-Rosenblum theorem “fiberwise” to deduce the first assertion. See Section 4.5.1 for more details, in particular Theorem 4.22. Assertion (ii) is a direct consequence of Assertion (i) together with the theory of direct integrals.
Theorem 3.11 refers to the unbound pair (scattering) channel. The subspace $\mathfrak{H}_{f}\subseteq\mathfrak{H}$ corresponds to the “incoming” ($+$) and “outcoming” (--) scattering states of the quantum system, in this particular scattering channel. Physically, this theorem shows, among other things, that the bosonic component of eitH vanishes on this channel, as $t\rightarrow\pm\infty$.
@@
$W^{\pm}\left(H,U_{f}^{\ast}\mathrm{H}_{f}U_{f};U^{\ast}\mathfrak{U}U_{f}\right)=U^{\ast}W^{\pm}\left(UHU^{\ast},\mathrm{H}_{f};\mathfrak{U}\right)U_{f}\ ,$
where $U_{f}$ is defined by (25). See also Equation (24), Proposition 2.1 and Section 5.1.
Note that, in any given fiber, the wave operators, as well as the scattering operators, can be computed even more explicitly via infinite series (perturbative expansions), thanks to Corollary 5.2. Below we give an example of such a computation by taking $\mathrm{U}=\mathrm{V}\in\mathbb{R}_{0}^{+}$ and $\mathrm{v}=\mathrm{u}:\mathbb{Z}^{2}\rightarrow\mathbb{R}_{0}^{+}$ in Theorem 3.11. With this particular choice, the unbound pair channel allows one to isolate the fermion-boson exchange mechanism, which in terms of Hamiltonians refers to the use of off-diagonal operators
@@
$T$ $<$ $t\implies\left\|\left(W^{+}\left(UHU^{\ast},\mathrm{H}_{f};\mathfrak{U}\right)-V_{0,t}\mathfrak{U}\right)\varphi\right\|_{\mathcal{X}}\leq\varepsilon\ ,$
$t$ $<$ $-T\implies\left\|\left(W^{-}\left(UHU^{\ast},\mathrm{H}_{f};\mathfrak{U}\right)-V_{0,t}\mathfrak{U}\right)\varphi\right\|_{\mathcal{X}}\leq\varepsilon\ ,$
@@
$s<-T<T<t\implies\left\langle\psi,S\left(UHU^{\ast},\mathrm{H}_{f};\mathfrak{U}\right)\varphi\right\rangle_{\mathcal{X}}=\left\langle\mathfrak{U}\psi,V_{t,s}\mathfrak{U}\varphi\right\rangle_{\mathcal{X}}+\mathcal{O}\left(\varepsilon\right)\ ,$
@@
$B_{1,2}^{(t)}\left(k\right)B_{2,1}^{(s)}\left(k\right)$ $=$ $\left(\hat{\upsilon}\left(k\right)^{2}\mathrm{e}^{i\left(s-t\right)\mathfrak{b}\left(k\right)}\right)\mathrm{e}^{itA_{1,1}\left(\mathrm{U},k\right)}P_{\mathfrak{d}\left(k\right)}\mathrm{e}^{-isA_{1,1}\left(\mathrm{U},k\right)}\ ,$
$B_{2,1}^{(t)}\left(k\right)B_{1,2}^{(s)}\left(k\right)$ $=$ $\left(\hat{\upsilon}\left(k\right)^{2}\mathrm{e}^{i\left(t-s\right)\mathfrak{b}\left(k\right)}\right)\left\langle\mathfrak{d}\left(k\right),\mathrm{e}^{i(s-t)A_{1,1}\left(\mathrm{U},k\right)}\mathfrak{d}\left(k\right)\right\rangle\ ,$
@@ 3.3.2 Bound pair scattering channel
Similarly, we also prove the existence of a scattering channel for dressed bound pairs. As dressed bound pairs are space-localized objects (see, e.g., Theorem 3.1 (iii)), the fermions forming the pair efficiently exchange a boson, at a non-negligible rate, via the terms $W_{\mathrm{b\rightarrow f}}$ (9) and $W_{\mathrm{f\rightarrow b}}$ (15) in the Hamiltonian $H$. In particular, such quantum states must have some non-negligible bosonic component representing the exchanged boson that “glues” the two fermions together. This dressed bound pair is however expected to move like a free (quantum spinless) particle in the real space. We translate this physical heuristics in precise mathematical terms by considering the effective dispersion relations
@@
    $\mathrm{e}^{itUHU^{\ast}}\mathfrak{P}_{\mathrm{U}}=\mathfrak{P}_{\mathrm{U}}\mathrm{e}^{itM_{\mathrm{E}\left(\mathrm{U},\cdot\right)}}\ ,\qquad t\in\mathbb{R}\ .$
@@
    $s-{\lim\limits_{\mathrm{U}\rightarrow\infty}}\mathfrak{P}_{\mathrm{U}}=\mathfrak{P}_{\infty}\qquad\text{and}\qquad s-{\lim\limits_{\mathrm{U}\rightarrow\infty}}\,\mathrm{e}^{itUHU^{\ast}}\mathfrak{P}_{\mathrm{U}}=\mathfrak{P}_{\infty}\mathrm{e}^{itM_{\mathrm{E}\left(\mathrm{\infty},\cdot\right)}}\ ,\qquad t\in\mathbb{R}\ .$
@@
$W^{\pm}\left(UHU^{\ast},M_{\mathrm{E}\left(\mathrm{U},\cdot\right)};\mathfrak{P}_{\mathrm{U}}\right)=\mathfrak{P}_{\mathrm{U}}P_{\mathrm{a}\mathrm{c}}\left(M_{\mathrm{E}\left(\mathrm{U},\cdot\right)}\right)$
@@
$S\left(UHU^{\ast},M_{\mathrm{E}\left(\mathrm{U},\cdot\right)};\mathfrak{P}_{\mathrm{U}}\right)=P_{\mathrm{a}\mathrm{c}}\left(M_{\mathrm{E}\left(\mathrm{U},\cdot\right)}\right)\ .$
@@
$W^{\pm}\left(UHU^{\ast},M_{\mathrm{E}\left(\mathrm{U},\cdot\right)};\mathfrak{P}_{\mathrm{U}}\right)=\mathfrak{P}_{\mathrm{U}}\qquad\text{and}\qquad S\left(UHU^{\ast},M_{\mathrm{E}\left(\mathrm{U},\cdot\right)};\mathfrak{P}_{\mathrm{U}}\right)=\mathfrak{1}$
@@
$\mathcal{P}_{\mathrm{U}}\doteq U^{\ast}\mathfrak{P}_{\mathrm{U}}U_{f}\ ,\qquad\mathrm{U}\in\mathbb{R}_{0}^{+}\ ,$
and $U_{f}\doteq U_{2}U_{1}$. See Equations (24)–(25), Proposition 2.1 and Section 5.1.

### v2

@@ 3.3.1 Unbound pair scattering channel
Far apart from each other, two fermions only experience a very weak repulsion force due to the extended Hubbard interaction while the probability that they bind together to form a boson is also very small. Thus, in this situation, one expects that the dynamics of such a pair is governed by the fermionic part, and even by the hopping term alone. During intermediate times they could of course interact, as they may get close to each other, and they could even be bound together via the effective attraction caused by fermion-boson exchange processes. The lifetime of bound fermions should however be finite in this situation and they are expected to be released at some point and behave again as two free fermions that go far apart from each other for large times. See Fig. 2. We show below that this heuristics can be put in precise mathematical terms.
[…]
Figure 2: Illustration of the unbound pair scattering channel: Two free fermions of (quasi-) momentum $k-p$ and $q$ respectively (i.e. the full momentum of the fermionic pair is $k$) at time $t=-\infty$ interact in finite time with the composite system, in particular with the bosonic field, to be asymptotically free again at time $t=+\infty$, thanks to Theorem 3.11. Here $S_{k}=S\left(A\left(k\right),\left(M_{\mathfrak{f}\left(k\right)}+R\left(\mathrm{V},\mathrm{v}\right)\right)\oplus A_{2,2}\left(k\right)\right)$ is the scattering operator of this process in each fiber $k$, which depends explicitly on $\hat{\upsilon}\left(k\right)$. See Theorem 3.13 and the example given by Equations (67)–(68).
@@
$M_{\mathfrak{f}\left(k\right)}$ being the fiber Hamiltonian defined as the multiplication operator by $\mathfrak{f}(k)\in C(\mathbb{T}^{2})$ (see (33)) while $P_{x}$ is the orthogonal projection onto the one-dimensional subspace $\mathbb{C}\mathfrak{\hat{e}}_{x}\subseteq L^{2}(\mathbb{T}^{2})$. Observe then that
$\mathbb{U}H\mathbb{U}^{\ast}-\begin{pmatrix}\mathrm{H}_{f}&0\\[5.0pt] 0&0\end{pmatrix}={\int_{\mathbb{T}^{2}}^{\oplus}}\begin{pmatrix}{\sum\limits_{x\in\mathbb{Z}^{2}}}\,\left(\mathrm{u}\left(x\right)-\mathrm{v}\left(x\right)\right)P_{x}+\left(\mathrm{U}-\mathrm{V}\right)P_{0}&A_{1,2}\left(k\right)\\[5.0pt] A_{2,1}\left(k\right)&A_{2,2}\left(k\right)\end{pmatrix}\,\nu(\mathrm{d}k)\ .$
Compare indeed (58) with Equations (46)–(37) and Proposition 2.1. By Lemma 4.21, note that $P_{\mathrm{ac}}(\mathrm{H}_{f})=\mathfrak{1}$.
@@
See Equation (25). Note that $\mathfrak{U}$ is an isometry, i.e., a norm preserving linear transformation. In fact, it is the canonical fiberwise inclusion of $\mathfrak{H}_{f}$ into $L^{2}\left(\mathbb{T}^{2},\mathcal{H}\right)$. Recall from Proposition 2.1 that $A\left(\mathrm{U},\cdot\right)$, defined by (37)–(46), is the fiber decomposition of the operator $\mathbb{U}H\mathbb{U}^{\ast}$. Then, we obtain wave and scattering operators with respect to fermionic parts:
@@
    The wave operators, as defined by (54) for $X=\mathbb{U}H\mathbb{U}^{\ast}$, $Y=\mathrm{H}_{f}$ and $J=\mathfrak{U}$, satisfy
    $W^{\pm}\left(\mathbb{U}H\mathbb{U}^{\ast},\mathrm{H}_{f};\mathfrak{U}\right)=\left({\int_{\mathbb{T}^{2}}^{\oplus}}W^{\pm}\left(A\left(k\right),\left(M_{\mathfrak{f}\left(k\right)}+R\left(\mathrm{V},\mathrm{v}\right)\right)\oplus A_{2,2}\left(k\right)\right)\nu\left(\mathrm{d}k\right)\right)\mathfrak{U}$ (61)
@@
    $\mathrm{ran}\left(W^{\pm}\left(\mathbb{U}H\mathbb{U}^{\ast},\mathrm{H}_{f};\mathfrak{U}\right)\right)=\int_{\mathbb{T}^{2}}^{\oplus}L^{2}\left(\mathbb{T}^{2}\right)\oplus\{0\}\,\nu\left(\mathrm{d}k\right)\ .$ (62)
@@
    The scattering operator, as defined by (56) for $X=\mathbb{U}H\mathbb{U}^{\ast}$, $Y=\mathrm{H}_{f}$ and $J=\mathfrak{U}$, equals
    $S\left(\mathbb{U}H\mathbb{U}^{\ast},\mathrm{H}_{f};\mathfrak{U}\right)=\mathfrak{U}^{\ast}\left({\int_{\mathbb{T}^{2}}^{\oplus}}S\left(A\left(k\right),\left(M_{\mathfrak{f}\left(k\right)}+R\left(\mathrm{V},\mathrm{v}\right)\right)\oplus A_{2,2}\left(k\right)\right)\nu\left(\mathrm{d}k\right)\right)\mathfrak{U}\ .$
Proof. Observe that the operator difference $(\mathbb{U}H\mathbb{U}^{\ast}-\mathrm{H}_{f})$ is not trace-class (it is not even compact) and, thus, the existence of this scattering channel is not a direct consequence of the well-known Kato-Rosenblum theorem [41, Theorem XI.8]. In fact, one of the main steps of the proof is to show that this difference is the direct integral of a strongly measurable family of trace-class operators. By this means we are then able to apply the Kato-Rosenblum theorem “fiberwise” to deduce the first assertion. See Section 4.6.1 for more details, in particular Theorem 4.23. Assertion (ii) is a direct consequence of Assertion (i) together with the theory of direct integrals.
@@
$W^{\pm}\left(H,U_{f}^{\ast}\mathrm{H}_{f}U_{f};\mathbb{U}^{\ast}\mathfrak{U}U_{f}\right)=U^{\ast}W^{\pm}\left(\mathbb{U}H\mathbb{U}^{\ast},\mathrm{H}_{f};\mathfrak{U}\right)U_{f}\ ,$
where $U_{f}$ is defined by (28). See also Equation (26) and Proposition 2.1.
Theorem 3.11 refers to the unbound pair (scattering) channel. The subspace $\mathfrak{H}_{f}\subseteq\mathfrak{H}$ corresponds to the “incoming” ($+$) and “outcoming” (--) scattering states of the quantum system, in this particular scattering channel. Physically, this theorem shows, among other things, that the bosonic component of eitH vanishes on this channel, as $t\rightarrow\pm\infty$. This is a direct consequence of Equation (62).
In addition, Equation (61) gives an explicit fiber decomposition of wave operators with respect to the purely fermionic Hamiltonian in terms of $k$-dependent wave operators defined naturally from the fiber decomposition of the operator $\mathbb{U}H\mathbb{U}^{\ast}$. Mutatis mutandis for the scattering operator, thanks to Theorem 3.11 (ii). In other words, the knowledge of scattering properties of each fiber, almost everywhere, entirely determines the scattering properties of the composite system, made of two fermions and one boson. We can now use this property, i.e. Theorem 3.11, to obtain a more computable expression for the wave and scattering operators in each given fiber. This can be done via infinite series (perturbative expansions), thanks to Corollary 5.2.
Below we give an example of such a computation by taking $\mathrm{U}=\mathrm{V}\in\mathbb{R}_{0}^{+}$ and $\mathrm{v}=\mathrm{u}:\mathbb{Z}^{2}\rightarrow\mathbb{R}_{0}^{+}$ in Theorem 3.11. With this particular choice, the unbound pair channel allows one to isolate the fermion-boson exchange mechanism, which in terms of Hamiltonians refers to the use of off-diagonal operators
@@
$T$ $<$ $t\implies\left\|\left(W^{+}\left(\mathbb{U}H\mathbb{U}^{\ast},\mathrm{H}_{f};\mathfrak{U}\right)-V_{0,t}\mathfrak{U}\right)\varphi\right\|_{\mathcal{X}}\leq\varepsilon\ ,$
$t$ $<$ $-T\implies\left\|\left(W^{-}\left(\mathbb{U}H\mathbb{U}^{\ast},\mathrm{H}_{f};\mathfrak{U}\right)-V_{0,t}\mathfrak{U}\right)\varphi\right\|_{\mathcal{X}}\leq\varepsilon\ ,$
@@
$s<-T<T<t\implies\left\langle\psi,S\left(\mathbb{U}H\mathbb{U}^{\ast},\mathrm{H}_{f};\mathfrak{U}\right)\varphi\right\rangle_{\mathcal{X}}=\left\langle\mathfrak{U}\psi,V_{t,s}\mathfrak{U}\varphi\right\rangle_{\mathcal{X}}+\mathcal{O}\left(\varepsilon\right)\ ,$
@@
$B_{1,2}^{(t)}\left(k\right)B_{2,1}^{(s)}\left(k\right)$ $=$ $\left(\hat{\upsilon}\left(k\right)^{2}\mathrm{e}^{i\left(s-t\right)\mathfrak{b}\left(k\right)}\right)\mathrm{e}^{itA_{1,1}\left(\mathrm{U},k\right)}P_{\mathfrak{d}\left(k\right)}\mathrm{e}^{-isA_{1,1}\left(\mathrm{U},k\right)}\ ,$ (67)
$B_{2,1}^{(t)}\left(k\right)B_{1,2}^{(s)}\left(k\right)$ $=$ $\left(\hat{\upsilon}\left(k\right)^{2}\mathrm{e}^{i\left(t-s\right)\mathfrak{b}\left(k\right)}\right)\left\langle\mathfrak{d}\left(k\right),\mathrm{e}^{i(s-t)A_{1,1}\left(\mathrm{U},k\right)}\mathfrak{d}\left(k\right)\right\rangle\ ,$ (68)
@@
The understanding of this kind of scattering, regarding free fermion (electron) collisions, is relevant in physics, because it can allow the exchange function $\upsilon$ of a real system to be studied. It is therefore important to have a model from which not only qualitative, but also quantitative, information can be obtained. This is the purpose of this section, in particular of Theorem 3.13, which gives the explicit dependency of scattering in terms of $\upsilon$.
@@ 3.3.2 Bound pair scattering channel
Similarly, we also prove the existence of a scattering channel for dressed bound pairs. As dressed bound pairs are space-localized objects (see, e.g., Theorem 3.1 (iii)), the fermions forming the pair efficiently exchange a boson, at a non-negligible rate, via the terms $W_{\mathrm{b\rightarrow f}}$ (9) and $W_{\mathrm{f\rightarrow b}}$ (17) in the Hamiltonian $H$. In particular, such quantum states must have some non-negligible bosonic component representing the exchanged boson that “glues” the two fermions together. This dressed bound pair is however expected to move like a free (quantum spinless) particle in the real space. See Fig. 3. We translate this physical heuristics in precise mathematical terms by considering the effective dispersion relations
@@
[…]
Figure 3: Illustration of the bound pair scattering channel. Here $k$ is the full (quasi-)momentum of the (exponentially localized) dressed bound fermion pairs. The oscillating vertical lines between the two fermions (e.g. electrons) before the scattering process and afterwards characterise their bound via a bosonic (e.g., bipolaronic) particle transfer with coupling function $\hat{\upsilon}\left(k\right)$, see Fig. 1. It illustrates the stability of these pairs of fermions in time, as expressed by Theorem 3.14, i.e., the pairs cannot decay into an (even only asymptotically) unbound pair of fermions.
@@
    $\mathrm{e}^{it\mathbb{U}H\mathbb{U}^{\ast}}\mathfrak{P}_{\mathrm{U}}=\mathfrak{P}_{\mathrm{U}}\mathrm{e}^{itM_{\mathrm{E}\left(\mathrm{U},\cdot\right)}}\ ,\qquad t\in\mathbb{R}\ .$
@@
    $s-{\lim\limits_{\mathrm{U}\rightarrow\infty}}\mathfrak{P}_{\mathrm{U}}=\mathfrak{P}_{\infty}\qquad\text{and}\qquad s-{\lim\limits_{\mathrm{U}\rightarrow\infty}}\,\mathrm{e}^{it\mathbb{U}H\mathbb{U}^{\ast}}\mathfrak{P}_{\mathrm{U}}=\mathfrak{P}_{\infty}\mathrm{e}^{itM_{\mathrm{E}\left(\mathrm{\infty},\cdot\right)}}\ ,\qquad t\in\mathbb{R}\ .$
@@
$W^{\pm}\left(\mathbb{U}H\mathbb{U}^{\ast},M_{\mathrm{E}\left(\mathrm{U},\cdot\right)};\mathfrak{P}_{\mathrm{U}}\right)=\mathfrak{P}_{\mathrm{U}}P_{\mathrm{a}\mathrm{c}}\left(M_{\mathrm{E}\left(\mathrm{U},\cdot\right)}\right)$
@@
$S\left(\mathbb{U}H\mathbb{U}^{\ast},M_{\mathrm{E}\left(\mathrm{U},\cdot\right)};\mathfrak{P}_{\mathrm{U}}\right)=P_{\mathrm{a}\mathrm{c}}\left(M_{\mathrm{E}\left(\mathrm{U},\cdot\right)}\right)\ .$
@@
$W^{\pm}\left(\mathbb{U}H\mathbb{U}^{\ast},M_{\mathrm{E}\left(\mathrm{U},\cdot\right)};\mathfrak{P}_{\mathrm{U}}\right)=\mathfrak{P}_{\mathrm{U}}\qquad\text{and}\qquad S\left(\mathbb{U}H\mathbb{U}^{\ast},M_{\mathrm{E}\left(\mathrm{U},\cdot\right)};\mathfrak{P}_{\mathrm{U}}\right)=\mathfrak{1}$
@@
$\mathcal{P}_{\mathrm{U}}\doteq\mathbb{U}^{\ast}\mathfrak{P}_{\mathrm{U}}U_{f}\ ,\qquad\mathrm{U}\in\mathbb{R}_{0}^{+}\ ,$
and $U_{f}\doteq U_{2}U_{1}$. See Equations (26)–(28) and Proposition 2.1.